OpenStudy (anonymous):

A triangle ABC is inscribed between the graphs of a sine curve and a cosine curve. f(x) = 5 sin x g(x) = 5 cos x If the period of both curves is 2Pi, find the area of triangle ABC, given point C is at the maxima of f(x) and points A and B both touch g(x), and f(x) respectively where g(x) = f(x) = 0.

OpenStudy (anonymous):

i think this is the answer: f(x) = 5 sin x = 0 g(x) = 5 cos x = 0 x = sin-1(0) = 0, Pi, 2Pi x = cos-1(0) = Pi/2, Pi, 3Pi/2 (0,0) B= (Pi,0) (2Pi,0) A= (Pi/2,0) (3Pi/2, 0)

OpenStudy (anonymous):

 I don't understand where g(x)=f(x)=0